A car starts from rest and accelerates uniformly on a rectilinear road.It reaches a speed of 72 km/h after 10s calculate its acceleration
1
Expert's answer
2020-04-08T10:52:56-0400

First of all convert the given speed into m/s

72x"\frac{5}{18}" =20m/s

As the car starts from rest then its initial velocity will be 0 and final velocity is 20 m/s and all this happens in a time span of sec

applying v=u+at

20=0+a.10

a=2 m/s2